### v1

@@ 6.1   Testing for the CRS in a production function
In this subsection, we are interested in studying whether the production function in China’s nonmetal mineral manufacturing industry has a constant returns to scale (CRS) technology. CRS refers to a scenario where a simultaneous proportional increase in all inputs results in an identical proportional increase in output. Returns to scale is a classical concept in economics to assess the efficiency of a production function, and it has accordingly received much attention in the empirical literature (for recent works, see Ackerberg et al., 2015; Basu and Bundick, 2017; Attanasio et al., 2020; Combes et al., 2021). Existing works, however, mostly do not account for cross-sectional dependence between firms. The role of cross-firms correlation has recently been highlighted empirically by Iyoha (2023), who shows that this can capture spillover effects and would lead to model misspecification if left unaccounted for.
We consider models that generalize Li et al. (2002)’s production function by incorporating spatial dependence in the analysis, thereby allowing for cross-sectional dependence. Their analysis centers on a Cobb-Douglas style production function that allows the elasticities of capital and labor to be nonparametrically varying with management expenses. Management expenses are costs that are indirectly associated with output production, such as research and development (R&D), equipment upgrades and employee training. In this framework, CRS corresponds to the elasticities of capital and labor summing to one.
Using data from the Third Industrial Census of China (conducted by the National Statistical Bureau in 1995) on the non-metal mineral manufacturing industry, Li et al. (2002) find that ignoring the varying nature of these coefficients leads to an underestimation of the returns to scale. Moreover, they suggest their estimate of the returns to scale of the technology is close to being constant for most values of managerial expense, but they do not have a test for the CRS hypothesis.
Our data comes from the Chinese Industrial Enterprise Database for the year of 2014 as this is the most recent information available in the database. We use data on all firms in the nonmetal mineral manufacturing industry with the survey code 31, resulting in a sample of 7,355 observations. As advocated by Li et al. (2002), the appeal for focusing on this industry is that production functions of firms are expected to be homogeneous since it has a very small proportion of foreign entity ownership. This is important because evidence suggests that production performance may vary across ownership types (Murakami et al., 1994). Table 1 introduces all variables after taking logarithms.
Our variables are output ($y_{i}$), capital ($p_{i1}$), labor ($p_{i2}$), and management expenses ($z_{i}$). As in Li et al. (2002), we include a parametric benchmark where the Cobb-Douglas production function has a quadratic form of management expenses without the varying coefficient entering additively, namely:
@@
The elasticities of capital and labor are $\delta_{1}$ and $\delta_{2}$, respectively. The semiparametric production function where the elasticities of of capital and labor vary with management expenses is:
@@
Based on the CRS hypothesis, $H_{0}^{true}:\delta_{1}(z_{i})+\delta_{2}(z_{i})=1$, we reparameterize the model as
@@
where $y_{i}^{\ast}=y_{i}-p_{i1}$. Then, we estimate
$\begin{split}&(\mathbb{W}_{0})\ y_{i}^{*}=\beta_{0}+\sum\limits_{k=1}^{h}\alpha_{1k}\psi_{ik}+\sum\limits_{k=1}^{h}\alpha_{2k}(p_{i2}-p_{i1})\psi_{ik}+\sum\limits_{k=1}^{h}\alpha_{3k}p_{i1}\psi_{ik}+\epsilon_{i},\end{split}$ (6.2)
where $\psi_{ik}$ is a basis function, with argument $z_{i}$, for $k=1,...,h/2$. We approximate the unknown functions by polynomial and trigonometric basis functions of different orders. We define the series approximation based Wald statistic derived from this model with no spatial autoregressive component as $\mathbb{W}_{0}$.
Next, we incorporate the SAR structure and consider:
@@
$\begin{split}&(\mathbb{W}_{1})\ y_{i}^{*}=\beta_{0}+\lambda\sum\limits_{j=1}^{n}w_{ij}y_{j}+\sum\limits_{k=1}^{h}\alpha_{1k}\psi_{ik}+\sum\limits_{k=1}^{h}\alpha_{2k}(p_{i2}-p_{i1})\psi_{ik}+\sum\limits_{k=1}^{h}\alpha_{3k}p_{i1}\psi_{ik}+\epsilon_{i},\\ &(\mathbb{W}_{2})\ y_{i}^{*}=\beta_{0}+\lambda\sum\limits_{j=1}^{n}w_{ij}y_{j}+\sum\limits_{k=1}^{h}\alpha_{1k}\psi_{ik}+\sum\limits_{k=1}^{h}\alpha_{2k}(p_{i2}-p_{i1})\psi_{ik}+\sum\limits_{k=1}^{h}\alpha_{3k}p_{i1}\psi_{ik}+\epsilon_{i},\\ &(\mathbb{W}_{3})\ y_{i}^{*}=\beta_{0}+\sum\limits_{l=1}^{h}\tau_{l}c_{il}+\sum\limits_{k=1}^{h}\alpha_{1k}\psi_{ik}+\sum\limits_{k=1}^{h}\alpha_{2k}(p_{i2}-p_{i1})\psi_{ik}+\sum\limits_{k=1}^{h}\alpha_{3k}p_{i1}\psi_{ik}+\epsilon_{i},\\ &(\mathbb{W}_{4})\ y_{i}^{*}=\beta_{0}+\sum\limits_{l=1}^{h}\mu_{l}h_{il}+\sum\limits_{k=1}^{h}\alpha_{1k}\psi_{ik}+\sum\limits_{k=1}^{h}\alpha_{2k}(p_{i2}-p_{i1})\psi_{ik}+\sum\limits_{k=1}^{h}\alpha_{3k}p_{i1}\psi_{ik}+\epsilon_{i},\end{split}$ (6.3)
@@
with $\epsilon_{i}=\sum\limits_{j=1}^{n}b_{ij}v_{j}$ for $\mathbb{W}_{2}$, $\mathbb{W}_{3}$ and $\mathbb{W}_{4}$. We use two different row-normalized spatial weight matrices, denoted $W_{n}^{p}$ and $W_{n}^{d}$:
@@
$\begin{split}&w_{ij}=[W_{n}^{p}]_{ij}=\begin{cases}1\quad\text{iffirms}\ i\ \text{and}\ j\ \text{are in the same prefecture-level city},\\ 0\quad\text{otherwise},\end{cases}\\ &w_{ij}=[W_{n}^{d}]_{ij}=\begin{cases}\frac{1}{d_{ij,m}^{*}}\quad\text{if}\ d_{ij,m}^{*}<d_{m},\\ 0\quad\text{otherwise},\end{cases}\end{split}$
where $d_{ij,m}^{\ast}$ measures the geographical distance between firms $i$ and $j$, and $d_{m}$ represents the 10th percentile of all geographical distances.
@@
For $\mathbb{W}_{i}$, $i=0,...,3$, $\alpha_{3}=(\alpha_{31},...\alpha_{3h})^{\prime}$, and the null is $H_{0}:\alpha_{3}=0$. For $\mathbb{W}_{4}$, the null is $H_{0}:\alpha_{3}=0$ or $H_{0}:\mu=0$, where $\mu=(\mu_{1},...,\mu_{h})^{\prime}$. Throughout our empirical analysis, we use $\psi_{ik}(z)=z_{i}^{k}$ as the polynomial function for $k=1,...,h$, and $\psi_{ik}(z)=[\sin(kz_{i}),\cos(kz_{i})]$ as the trigonometric function $k=1,...,h/2$ for $h=2,4$. To compute $\mathbb{W}_{2}$, $\mathbb{W}_{3}$ and $\mathbb{W}_{4}$, we use the Epanechnikov kernel function to estimate $\mathscr{K}\left({d_{ij,m}^{\ast}}/{d_{m}}\right)$. For $\mathbb{W}_{3}$, we denote $\widehat{w}_{ij,l}=[d_{ij,m}^{\ast}]^{l}\mathbf{1}(d_{ij,m}^{\ast}<d_{m})$ if $i\neq j$ for $l=1,...,h$. Then, $c_{il}=\sum_{j=1}^{n}\widehat{w}_{ij,l}y_{j}$ for $l=1,...,h$. For $\mathbb{W}_{4}$, we use $\phi_{il}(z)=[\frac{2}{\pi}\tanh(z_{i})]^{l}$ as the polynomial function and $\phi_{il}(z)=\sin(\frac{z_{i}}{2l})$ as the trigonometric function for $l=1,...,h$. Then, $h_{il}=y_{i}\sum_{j=1}^{n}w_{ij}\phi_{jl}(z)$. For the IVs, we use $K_{1}=[W\Psi,X,\Psi]$ with respect to $\mathbb{W}_{1}$ and $\mathbb{W}_{2}$, and $K_{2}=[X,\Psi]$ and $K_{3}=[H\Psi,X,\Psi]$ with respect to $\mathbb{W}_{3}$ and $\mathbb{W}_{4}$.
The results from estimating the parametric model can be found in Table 2, and the results for other models with different basis functions are in Tables 3-6. In the latter group of tables, we define $\mathscr{T}_{n}$ as the asymptotic p-values corresponding to $\mathbb{W}_{i}$, $i=0,...,4$. It is well known that nonparametric tests, which rely on their asymptotic normal null distributions, may not perform well with finite samples. As an alternative, we provide bootstrap p-values that are denoted by $\mathscr{T}_{n}^{\ast}$ basing on 150 bootstrap replications. The details of the bootstrap procedure are given in Section 7.1.
@@
Our results show that, in all models, the hypothesis of CRS cannot be rejected. These findings corroborate those in Li et al. (2002). A conclusion one can draw from this is that CRS is a salient feature of the production function in this industry, as it is present in the parametric model as well as in semiparametric models with varying coefficients and cross-firm dependence.
@@
We end this application with some remarks that, in addition to inference on returns to scale, our analysis also reveals other similarities with prior works. Notably, from Table 7, we observe that the parametric method underestimates the returns to scale compared to the semiparametric method. [^1] The returns to scale in a semiparametric model can be computed using $\frac{1}{n}\sum_{i=1}^{n}\hat{\delta}_{1}(z_{i})$ and $\frac{1}{n}\sum_{i=1}^{n}\hat{\delta}_{2}(z_{i})$. The returns to scale with respect to different input factors also differ, which highlights the crucial role of management expenses. To illustrate this, take the semiparametric estimates with spatial weights $W_{n}^{d}$ and produce Figure 1 and Figure 2. The former shows that the output elasticity of capital is increasing in management expenses, and the latter shows the output elasticity of labor is decreasing in management expenses.
Intuitively, by the nature of management expenses in production, the elasticities may be expected to increase with both inputs. However, the decreasing elasticities of labor is a known empirical fact in this industry due to the practice of “concealed unemployment”  where the government did not allow state-owned firms to lay off extra employees as a strategy to avoid social unrest. We also note that the returns to scale is relatively flat in the semiparametric model and is within the 95% bounds for most $z_{i}$, as illustrated in Figure 3. Further discussions on these findings can be found in Li et al. (2002). The results we have described here are not unique to the estimates from $W_{n}^{d}$. Other choices of spatial weights give the same conclusion; we simply selected one to illustrate our findings.
6.2   Testing the impact of the distance to the employment center on housing price
@@
In this subsection, we use the Boston house price dataset, which has been used in several spatial studies, to illustrate another application of our proposed test. The dataset [^2] The related shp file can be downloaded at https://github.com/simonbrewer/geog6000/blob/main/datafiles/boston.tr.zip. Data descriptions are in Table 8. consists of the median value of owner-occupied homes in 506 census tracts in the Boston Standard Metropolitan Statistical Area in 1970 (Harrison Jr and Rubinfeld, 1978). Specifically, we explore how some factors such as the per capita crime rate by town (denoted by CRIM), average number of rooms per dwelling (RM), full-value property-tax rate per \$10,000 dollar (TAX), percentage of lower socioeconomic status population (LSTAT) and index of accessibility to radial highways (RAD) affect the median value of owner-occupied homes in $1000$s (MEDV), and whether the effects of these factors vary over location or not. DIS, which represents the weighted distances between a property and the five Boston employment centers that are associated with higher housing prices due to closer proximity to the city center, will drive the varying coefficient.
@@
The motivation for considering spatial autocorrelation in house prices is straightforward, as housing prices in a specific location or region are commonly influenced by the prices of nearby houses due to factors such as environmental and geographic considerations. [^3] There are also some studies that use the SAR semiparametric model with the Boston house price data, for example, Malikov and Sun (2017) and Li et al. (2019). However, the former considers the covariate $z_{i}$ as NO2, which is more about the impact of environmental pollution on housing prices. The latter concentrates on the geographical weighted regression (GWR) method. Sun et al. (2014) proposed a semiparametric coefficient varying spatial model and derived the distribution theory of their estimator. While they do not have inference results on functions of the random coefficients, they have a model selection procedure to determine the parametric and nonparametric components in their model. We thus take the model they selected as our starting point.
@@
Let $y_{i},x_{i1},x_{i2},p_{i1},p_{i2},p_{i3},z_{i}$ denote the log of MEDV, RAD, LSTAT, CRIM, RM, TAX and DIS respectively. Guided by model selection results, Sun et al. (2014) argue that DIS doesn’t affect LSTAT and RAD. Intuitively, for LSTAT, the geographical distribution of employment opportunities does not necessarily correlate with a community’s socio-economic status; a community might be far from employment centers but could have good public transportation links or a high number of remote work opportunities, mitigating the impact on the community’s lower socioeconomic status. For RAD, the accessibility to highways might depend more on urban planning and infrastructure investment rather than the proximity to employment centers.
The baseline model, based on Sun et al. (2014), is:
@@
We focus on a more general version of their model, by allowing the spatial lag coefficient to be varying with DIS as well:
@@
$\begin{split}&y_{i}=\lambda\sum\limits_{j=1}^{n}w_{ij}y_{j}+\sum\limits_{k=0}^{2}\beta_{k}x_{ik}+\sum\limits_{m=1}^{3}\delta_{m}(z_{i})p_{im}+\epsilon_{i},\ \text{and using}\ \mathbb{W}_{1};\\ &y_{i}=\lambda\sum\limits_{j=1}^{n}w_{ij}y_{j}+\sum\limits_{k=0}^{2}\beta_{k}x_{ik}+\sum\limits_{m=1}^{3}\delta_{m}(z_{i})p_{im}+\epsilon_{i},\ \text{with}\ \epsilon_{i}=\sum\limits_{j=1}^{n}b_{ij}v_{j},\ \text{and using}\ \mathbb{W}_{2}\\ &y_{i}=\sum\limits_{j=1}^{n}g(d_{ij})y_{j}+\sum\limits_{k=0}^{2}\beta_{k}x_{ik}+\sum\limits_{m=1}^{3}\delta_{m}(z_{i})p_{im}+\epsilon_{i},\ \text{with}\ \epsilon_{i}=\sum\limits_{j=1}^{n}b_{ij}v_{j},\ \text{and using}\ \mathbb{W}_{3};\\ &y_{i}=\lambda(z_{i})\sum\limits_{j=1}^{n}w_{ij}y_{j}+\sum\limits_{k=0}^{2}\beta_{k}x_{ik}+\sum\limits_{m=1}^{3}\delta_{m}(z_{i})p_{im}+\epsilon_{i},\ \text{with}\ \epsilon_{i}=\sum\limits_{j=1}^{n}b_{ij}v_{j},\ \text{and using}\ \mathbb{W}_{4},\end{split}$
and the estimated models are
$\begin{split}&y_{i}=\lambda\sum\limits_{j=1}^{n}w_{ij}y_{j}+\sum\limits_{k=0}^{2}\beta_{k}x_{ik}+\sum\limits_{m=1}^{3}\sum\limits_{k=1}^{h}\alpha_{mk}p_{im}\psi_{ik}+\epsilon_{i},\ \text{and using}\ \mathbb{W}_{1};\\ &y_{i}=\lambda\sum\limits_{j=1}^{n}w_{ij}y_{j}+\sum\limits_{k=0}^{2}\beta_{k}x_{ik}+\sum\limits_{m=1}^{3}\sum\limits_{k=1}^{h}\alpha_{mk}p_{im}\psi_{ik}+\epsilon_{i},\ \text{with}\ \epsilon_{i}=\sum\limits_{j=1}^{n}b_{ij}v_{j},\ \text{and using}\ \mathbb{W}_{2};\\ &y_{i}=\sum\limits_{l=1}^{h}\tau_{l}c_{il}+\sum\limits_{k=0}^{2}\beta_{k}x_{ik}+\sum\limits_{m=1}^{3}\sum\limits_{k=1}^{h}\alpha_{mk}p_{im}\psi_{ik}+\epsilon_{i},\ \text{with}\ \epsilon_{i}=\sum\limits_{j=1}^{n}b_{ij}v_{j},\ \text{and using}\ \mathbb{W}_{3};\\ &y_{i}=\sum\limits_{l=1}^{h}\mu_{l}h_{il}+\sum\limits_{k=0}^{2}\beta_{k}x_{ik}+\sum\limits_{m=1}^{3}\sum\limits_{k=1}^{h}\alpha_{mk}p_{im}\psi_{ik}+\epsilon_{i},\ \text{with}\ \epsilon_{i}=\sum\limits_{j=1}^{n}b_{ij}v_{j},\ \text{and using}\ \mathbb{W}_{4}.\end{split}$ (6.4)
We consider two row-normalized spatial weight matrices. The first is the first-order queen matrix, $W_{n}^{q}$, discussed in Malikov and Sun (2017). The second, denoted by $W_{n}^{t}$, indicates whether two locations are in the same tract. For $\mathbb{W}_{i}$, $i=0,...,3$, $\alpha=(\alpha_{11},...\alpha_{1h},...,\alpha_{3h})^{\prime}$, and the null is $H_{0}:\alpha=0$. For $\mathbb{W}_{4}$, the null is $H_{0}:\alpha=0$ or $H_{0}:\mu=0$, where $\mu=(\mu_{1},...,\mu_{h})^{\prime}$.
From Tables 9-12, we find that $log(DIS)$ indeed affects $p_{i1}$, $p_{i2}$ and $p_{i3}$, and has a significantly nonlinear impact on neighboring house prices. All test statistics show a rejection of the null $H_{0}:\alpha=0$. The bootstrap p-value also confirms this conclusion. In addition, testing with $\mathbb{W}_{4}$ also leads to a rejection of the null $H_{0}:\mu=0$, indicating a significant nonlinear relationship between $log(DIS)$ and neighboring house prices. Furthermore, we estimate the varying $\lambda(z_{i})$ and evaluate its empirical mean $\hat{\lambda}=\frac{1}{n}\sum_{i=1}^{n}\lambda(z_{i})$. From Table 13, we thus find that the parametric method underestimates the spatial coefficient. Our results, which are based on more general models, thus complement those in Sun et al. (2014).
7.   Monte Carlo Simulation
We present a basic set of Monte Carlo results in the main body of the paper. Because our theory covers such a wide range of models, which include several important special cases, we have performed a wide range of simulations for them too. However, the sheer number of tables we have makes it impractical to include them all in the main body of the paper, and we provide references to an online supplementary appendix.

### v2

@@ 6.1 Testing for the CRS in a production function
In this subsection, we are interested in testing whether the production function in China's nonmetal mineral manufacturing industry has a constant returns to scale (CRS) technology. CRS refers to a scenario where a simultaneous proportional increase in all inputs results in an identical proportional increase in output. Returns to scale is a classical concept in economics to assess the efficiency of a production function, and it has accordingly received much attention (see Ackerberg et al. (2015); Basu and Bundick (2017); Attanasio et al. (2020); Combes et al. (2021)). Existing works, however, mostly do not account for cross-sectional dependence between firms. The role of cross-firm correlation has recently been highlighted empirically by Iyoha (2023), who shows that this can capture spillover effects and would lead to model misspecification if left unaccounted for.
Our models generalize Li et al. (2002)'s production function by incorporating spatial dependence. Their analysis centers on a Cobb-Douglas production function that allows elasticities of capital and labor to be nonparametrically varying with management expenses. Management expenses are costs that are indirectly associated with output production, such as research and development (R&D), equipment upgrades and employee training. CRS corresponds to the elasticities of capital and labor summing to one.
Using data from the Third Industrial Census of China (conducted by the National Statistical Bureau in 1995), Li et al. (2002) find that ignoring the varying nature of these coefficients leads to an underestimation of the returns to scale. They suggest their estimate of returns to scale of the technology is close to being constant for most values of managerial expense, but they do not have a test for the CRS hypothesis. Our data comes from the Chinese Industrial Enterprise Database in 2014 as it is the most recent information available. We use all firms in the nonmetal mineral manufacturing industry with the survey code 31, resulting in a sample of 7,355 observations. As advocated by Li et al. (2002), production functions of firms in this industry are expected to be homogeneous since it has a very small proportion of foreign entity ownership. This is important because evidence suggests that production performance may vary across ownership types (Murakami et al. (1994)).
Variables are in logs: output ($y_{i}$), capital ($p_{i1}$), labor ($p_{i2}$), and management expenses ($z_{i}$). As in Li et al. (2002), we include a parametric benchmark model:
@@
Elasticities of capital and labor are $\delta_{1}$ and $\delta_{2}$, respectively. The semiparametric model where the elasticities of of capital and labor vary with management expenses is:
@@
Based on the CRS, $H_{0}^{true}:\delta_{1}(z_{i})+\delta_{2}(z_{i})=1$, we reparameterize the model as
@@
where $y_{i}^{\ast}=y_{i}-p_{i1}$. Then, we estimate the following model with no SAR structure:
$\begin{split}&(\mathbb{W}_{0})\ y_{i}^{*}=\beta_{0}+\sum\limits_{k=1}^{h}\alpha_{1k}\psi_{ik}+\sum\limits_{k=1}^{h}\alpha_{2k}(p_{i2}-p_{i1})\psi_{ik}+\sum\limits_{k=1}^{h}\alpha_{3k}p_{i1}\psi_{ik}+\epsilon_{i},\end{split}$
where $\psi_{ik}$ is a basis function, with argument $z_{i}$, for k=1,,h/2k=1,...,h/2. We approximate the unknown functions by polynomial and trigonometric functions of different orders. We define the series approximation based Wald statistic derived from this model as $\mathbb{W}_{0}$. Next, we incorporate the SAR structure and consider:
@@
$\begin{split}&(\mathbb{W}_{1})\ y_{i}^{*}=\beta_{0}+\lambda\sum\limits_{j=1}^{n}w_{ij}y_{j}+\sum\limits_{k=1}^{h}\alpha_{1k}\psi_{ik}+\sum\limits_{k=1}^{h}\alpha_{2k}(p_{i2}-p_{i1})\psi_{ik}+\sum\limits_{k=1}^{h}\alpha_{3k}p_{i1}\psi_{ik}+\epsilon_{i},\\ &(\mathbb{W}_{2})\ y_{i}^{*}=\beta_{0}+\lambda\sum\limits_{j=1}^{n}w_{ij}y_{j}+\sum\limits_{k=1}^{h}\alpha_{1k}\psi_{ik}+\sum\limits_{k=1}^{h}\alpha_{2k}(p_{i2}-p_{i1})\psi_{ik}+\sum\limits_{k=1}^{h}\alpha_{3k}p_{i1}\psi_{ik}+\epsilon_{i},\\ &(\mathbb{W}_{3})\ y_{i}^{*}=\beta_{0}+\sum\limits_{l=1}^{h}\tau_{l}c_{il}+\sum\limits_{k=1}^{h}\alpha_{1k}\psi_{ik}+\sum\limits_{k=1}^{h}\alpha_{2k}(p_{i2}-p_{i1})\psi_{ik}+\sum\limits_{k=1}^{h}\alpha_{3k}p_{i1}\psi_{ik}+\epsilon_{i},\\ &(\mathbb{W}_{4})\ y_{i}^{*}=\beta_{0}+\sum\limits_{l=1}^{h}\mu_{l}h_{il}+\sum\limits_{k=1}^{h}\alpha_{1k}\psi_{ik}+\sum\limits_{k=1}^{h}\alpha_{2k}(p_{i2}-p_{i1})\psi_{ik}+\sum\limits_{k=1}^{h}\alpha_{3k}p_{i1}\psi_{ik}+\epsilon_{i},\end{split}$
@@
with $\epsilon_{i}=\sum\limits_{j=1}^{n}b_{ij}v_{j}$ for $\mathbb{W}_{2}$, $\mathbb{W}_{3}$ and $\mathbb{W}_{4}$. Let $d_{ij,m}^{\ast}$ be the geographical distance between firms $i$ and $j$, and $d_{m}$ represent the 10th percentile of all geographical distances. We use two different row-normalized spatial weight matrices, denoted $W_{n}^{p}$ and $W_{n}^{d}$:
@@
$\begin{split}w_{ij}^{p}=\begin{cases}1\quad\text{if}\ i\ \text{and}\ j\ \text{are in the same city},\\ 0\quad\text{otherwise}.\end{cases}w_{ij}^{d}=\begin{cases}\frac{1}{d_{ij,m}^{*}}\quad\text{if}\ d_{ij,m}^{*}<d_{m},\\ 0\quad\text{otherwise}.\end{cases}\end{split}$
@@
For $\mathbb{W}_{i}$, i=0,,3i=0,...,3, $\alpha_{3}=(\alpha_{31},...\alpha_{3h})^{\prime}$, and the null is $H_{0}:\alpha_{3}=0$. For $\mathbb{W}_{4}$, the null is $H_{0}:\alpha_{3}=0$ or $H_{0}:\mu=0$, where $\mu=(\mu_{1},...,\mu_{h})^{\prime}$. Throughout our empirical analysis, we use $\psi_{ik}(z)=z_{i}^{k}$ as the polynomial function for k=1,,hk=1,...,h, and $\psi_{ik}(z)=[\sin(kz_{i}),\cos(kz_{i})]$ as the trigonometric function k=1,,h/2k=1,...,h/2 for $h=2,4$. To compute $\mathbb{W}_{2}$, $\mathbb{W}_{3}$ and $\mathbb{W}_{4}$, we use the Epanechnikov kernel function. For $\mathbb{W}_{3}$, let $\mathcal{E}_{l}$ be the matrix whose typical entry is, for $i\neq j$ and l=1,,hl=1,...,h, $e_{l}(d_{ij})=[d_{ij,m}^{\ast}]^{l}\mathbf{1}(d_{ij,m}^{\ast}<d_{m})$. Then, $c_{il}=\sum_{j=1}^{n}e_{l}(d_{ij})y_{j}$ for l=1,,hl=1,...,h. For $\mathbb{W}_{4}$, we use $\phi_{il}(z)=\frac{1}{h}[\frac{2}{\pi}\tanh(z_{i})]^{l}$ as the polynomial function and $\phi_{il}(z)=\frac{1}{h}\sin(\frac{z_{i}}{2l})$ as the trigonometric function for l=1,,hl=1,...,h. Then, $h_{il}=y_{i}\sum_{j=1}^{n}w_{ij}\phi_{jl}(z)$. For the IVs, we use $K_{1}=K_{3}=[X,\Psi,W_{n}^{p}X,W_{n}^{d}X]$ with respect to $\mathbb{W}_{1}$, $\mathbb{W}_{2}$ and $\mathbb{W}_{4}$, and $K_{2}=[X,\Psi,\mathcal{E}_{1}X,...,\mathcal{E}_{h}X]$ with respect to $\mathbb{W}_{3}$.
@@
Results with different basis functions are in Table 1. In all models, the hypothesis of CRS cannot be rejected. These findings corroborate those in Li et al. (2002). A conclusion one can draw from this is that CRS is a salient feature of the production function in this industry, as it is present in the parametric model as well as in semiparametric models with varying coefficients and cross-firm dependence.
@@
We end this application with some remarks: (1) In addition to inference on CRS, our analysis also reveals other similarities with prior works. Table 2 shows that the parametric method underestimates the returns to scale compared to the semiparametric method. [^1] The returns to scale in a semiparametric model is computed by $\frac{1}{n}\sum_{i=1}^{n}\hat{\delta}_{1}(z_{i})$ and $\frac{1}{n}\sum_{i=1}^{n}\hat{\delta}_{2}(z_{i})$. The returns to scale with respect to different input factors also differ, which highlights the crucial role of management expenses. Figures 1 (a)-(b) illustrates this with the semiparametric estimates with spatial weights $W_{n}^{d}$. They show that the output elasticity of capital is increasing in management expenses, while it is decreasing for labor. The decreasing elasticity of labor is a known empirical fact in this industry due to “concealed unemployment”  where the government did not allow state-owned firms to lay off extra employees as a strategy to avoid social unrest. (2) The returns to scale are relatively flat in the semiparametric model and are within the 95% bounds for most $z_{i}$, as illustrated in Figure 1 (c). Further discussions on these findings can be found in Li et al. (2002). (3) Our results are not unique to the estimates from $W_{n}^{d}$. Other choices of spatial weights give the same conclusion.
6.2 Testing the impact of the distance on house prices
@@
We use the Boston house price dataset  [^2] The dataset consists the median value of owner-occupied homes in 506 census tracts in the Boston Standard Metropolitan Statistical Area in 1970 (Harrison Jr and Rubinfeld (1978)), which is available at https://github.com/simonbrewer/geog6000/blob/main/datafiles/boston.tr.zip. to explore how some factors affect the median value of owner-occupied homes in $1000$s (MEDV), and whether the effects of these factors vary over location. Factors include the per capita crime rate by town (denoted by CRIM), average number of rooms per dwelling (RM), full-value property-tax rate per \$10,000 dollar (TAX), percentage of lower socioeconomic status population (LSTAT) and index of accessibility to radial highways (RAD). The location variable DIS, which represents the weighted distances between a property and the five Boston employment centers, will drive the varying coefficient. Let $y_{i},x_{i1},x_{i2},p_{i1},p_{i2},p_{i3},z_{i}$ denote the logarithm of MEDV, RAD, LSTAT, CRIM, RM, TAX and DIS respectively.
@@
As house prices in a specific location or region are commonly influenced by the prices of nearby houses due to factors such as environmental and geographic considerations, it is natural to consider the spatial autocorrelation. [^3] There are also some studies that use the SAR semiparametric model with the Boston house price data, for example, Malikov and Sun (2017) and Li et al. (2019). However, the former considers the covariate $z_{i}$ as NO2, which is more about the impact of environmental pollution on housing prices. The latter concentrates on the geographical weighted regression (GWR) method. Sun et al. (2014) proposed a semiparametric coefficient varying spatial model and derived the distribution theory of their estimator. While they do not have inference results on functions of the random coefficients, they have a model selection procedure to determine the parametric and nonparametric components. We thus take the model they selected as our starting point.
@@
Guided by model selection results, Sun et al. (2014) argue that DIS doesn't affect LSTAT and RAD. Intuitively, for LSTAT, the geographical distribution of employment opportunities does not necessarily correlate with a community's socio-economic status; a community might be far from employment centers but could have good public transportation links or a high number of remote work opportunities, mitigating the impact on the community's lower socioeconomic status. For RAD, the accessibility to highways might depend more on urban planning and infrastructure investment rather than the proximity to employment centers. The baseline model, based on Sun et al. (2014), is:
@@
We consider a more general version where the spatial lag coefficient varies with DIS:
@@
$\begin{split}&y_{i}=\lambda\sum\limits_{j=1}^{n}w_{ij}y_{j}+\sum\limits_{k=0}^{2}\beta_{k}x_{ik}+\sum\limits_{m=1}^{3}\sum\limits_{k=1}^{h}\alpha_{mk}p_{im}\psi_{ik}+\epsilon_{i},\ \text{and using}\ \mathbb{W}_{1};\\ &y_{i}=\lambda\sum\limits_{j=1}^{n}w_{ij}y_{j}+\sum\limits_{k=0}^{2}\beta_{k}x_{ik}+\sum\limits_{m=1}^{3}\sum\limits_{k=1}^{h}\alpha_{mk}p_{im}\psi_{ik}+\epsilon_{i},\ \ \epsilon_{i}=\sum\limits_{j=1}^{n}b_{ij}v_{j},\ \text{using}\ \mathbb{W}_{2};\\ &y_{i}=\sum\limits_{l=1}^{h}\tau_{l}c_{il}+\sum\limits_{k=0}^{2}\beta_{k}x_{ik}+\sum\limits_{m=1}^{3}\sum\limits_{k=1}^{h}\alpha_{mk}p_{im}\psi_{ik}+\epsilon_{i},\ \ \epsilon_{i}=\sum\limits_{j=1}^{n}b_{ij}v_{j},\ \text{using}\ \mathbb{W}_{3};\\ &y_{i}=\sum\limits_{l=1}^{h}\mu_{l}h_{il}+\sum\limits_{k=0}^{2}\beta_{k}x_{ik}+\sum\limits_{m=1}^{3}\sum\limits_{k=1}^{h}\alpha_{mk}p_{im}\psi_{ik}+\epsilon_{i},\ \ \epsilon_{i}=\sum\limits_{j=1}^{n}b_{ij}v_{j},\ \text{using}\ \mathbb{W}_{4}.\end{split}$
We consider two row-normalized spatial weight matrices. The first is the first-order queen matrix, $W_{n}^{q}$, discussed in Malikov and Sun (2017). The second, $W_{n}^{t}$, indicates whether two locations are in the same tract. For $\mathbb{W}_{i}$, i=0,,3i=0,...,3, $\alpha=(\alpha_{11},...\alpha_{1h},...,\alpha_{3h})^{\prime}$, and the null is $H_{0}:\alpha=0$. For $\mathbb{W}_{4}$, the null is $H_{0}:\alpha=0$ or $H_{0}:\mu=0$, where $\mu=(\mu_{1},...,\mu_{h})^{\prime}$.
Results are listed in Table 3. All tests reject the null $H_{0}:\alpha=0$. In addition, testing with $\mathbb{W}_{4}$ leads to a rejection of $H_{0}:\mu=0$, indicating a significant nonlinear relationship between $log(DIS)$ and neighboring house prices. We further estimate the varying $\lambda(z_{i})$ and evaluate its empirical mean $\hat{\lambda}=\frac{1}{n}\sum_{i=1}^{n}\lambda(z_{i})$. As shown in Panel E of Table 3, the parametric method underestimates the spatial coefficient. Thus our results based on more general models complement those in Sun et al. (2014).
7.  Monte Carlo
We present a basic set of Monte Carlo results in this section and a wider range of simulations in the online appendix. All experiments use 1000 replications.